OpenStudy (anonymous):

Multi-Step Problem A chauffeur can be paid in two ways. Plan A : $500.00 plus $9.00 per hour Plan B : $14.00 per hour Suppose that the job takes n hours. For what values of n is Plan A better for the chauffeur than Plan B

OpenStudy (theeric):

Hi! Any idea where you want to start?

OpenStudy (anonymous):

I need a starting equation for this problem, But I don't know how to set it up first.

OpenStudy (theeric):

Okay! I'll help you speak math! We'll say \(n\) is the number of hours a job can take. Then the cost of plan A, which I'll name \(C_A\) for cost of plan A, is \(C_A=$500+\dfrac{$9\ n}{[hr]}\quad\leftarrow\)$500 plus $9/hr times the number of hours the job takes. Do you want to take a shot at describing plan B?

OpenStudy (anonymous):

I'll try to ^^|| sorry if math isn't my best subject

OpenStudy (theeric):

It's not mine either, but I got through it a while ago and have practiced it as I've gotten into more math courses! I disliked math so strongly in high school. Now I see its use! I had a few good teachers to help me like it, but some other teachers ruined it for me. Take your time. And ask any questions you want. I think it's great that you're trying.

OpenStudy (anonymous):

It would probably be like this? Plan B , would start out as B for the result of the plan B = $14/{hr}

OpenStudy (theeric):

Well, the chauffeur is driving around and getting paid $14 each hour. So, if that chauffeur drove for, say, 5 hours, how much money would it cost?

OpenStudy (anonymous):

so that would mean he is getting $70 for 5 hours right?

OpenStudy (theeric):

Yeah! How did you get that?

OpenStudy (anonymous):

I multiplied 5 hours and the amount of money that he received and it equaled up to 70 dollars in the end

OpenStudy (theeric):

Right! So, if the chauffeur drives \(n\) hours, the total price for plan B will be $14/hr times \(n\), which looks like \(\dfrac{$14\ n}{[hr]}\). We'll call that \(C_B\), the cost of plan B. Does that sound good to you?

OpenStudy (anonymous):

yea

OpenStudy (theeric):

Alright! So now the problem wants to know when plan A is better than plan B. So, that would be when the cost of plan A is less than the cost of plan B. (It's better to pay less, right?) Have you worked with inequalities before? \(\le\ \ge\ <\ >\) stuff?

OpenStudy (anonymous):

yea, but I'm kind of getting the hang of those

OpenStudy (theeric):

That's good! :) Then we can use them. Like I said, we want to talk about when the cost of plan A is less than the cost of plan B. With my variables, that is \(C_A<C_B\). Is that okay so far?

OpenStudy (anonymous):

yea

OpenStudy (theeric):

Okay then. Now we look at what \(C_A\) and \(C_B\) really are. So \($500+\dfrac{$9\ n}{[hr]}<\dfrac{$14\ n}{[hr]}\) Now we can solve to see for what number of hours \((n)\) the cost of plan A is really less than plan B, which the inequality says for us. So, we should start solving for \(n\).

OpenStudy (theeric):

I would subtract \(\dfrac{$9\ n}{[hr]}\) from both sides to start. Sound good?

OpenStudy (anonymous):

yes

OpenStudy (theeric):

\($500<\dfrac{$14\ n}{[hr]}-\dfrac{$9\ n}{[hr]}=\dfrac{$5\ n}{[hr]}\) So \($500<\dfrac{$5\ n}{[hr]}\) Any questions? Double check me to make sure I didn't make a mistake!

OpenStudy (anonymous):

no, it's good

OpenStudy (theeric):

Okay, thanks. Now, \($500<\dfrac{$5\ n}{[hr]}\). To get \(n\) alone, what do you think we should do to both sides next?

OpenStudy (anonymous):

multiply the hr to both sides?

OpenStudy (theeric):

Okay! \($500\ [hr]<$5\ n\) And now what?

OpenStudy (anonymous):

divide 5 to both sides

OpenStudy (theeric):

Good plan! \(\dfrac{$500\ [hr]}{$5}=100\ [hr]<n\) The dollar units cancelled, and \(500\div 5=100\), so that's what I got. So, if the cost of plan A is less than plan B, then \(n\gt100\ [\text{hours}]\).

OpenStudy (theeric):

So your done! Any questions?

OpenStudy (anonymous):

nope, thanks for the help though ^^

OpenStudy (theeric):

You're very welcome! You did good to get through it, too :) Take care!

OpenStudy (anonymous):

you too :3

OpenStudy (theeric):

Thanks! :)

OpenStudy (anonymous):

your welcome